Circular No. 7369

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SAX J1808.4–3658 R. Wijnands, D. Chakrabarty, and E. Morgan, Massachusetts Insti-

tute of Technology; and M. van der Klis, University of Amsterdam, report: “Since the recent report of resumed activity from the transient millisecond x-ray pulsar SAX J1808.4–3658 (IAUC 7358), intermittent monitoring with the PCA instrument on the Rossi X-ray Timing Explorer has shown erratic variability of the source intensity on time scales shorter than a day. Between Feb. 8 and 25, the 2–10-keV x-ray intensity was sometimes as high as 12 mCrab but at other times undetected (< 1 mCrab). Detected emission during this interval clearly showed pulsations at the 401-Hz spin frequency but did not show the violent quasi-periodic x-ray flaring that was observed during Feb. 2–6 (IAUC 7358). The source was completely undetected during Feb. 20–25, but it has been intermittently detected since then at an intensity of roughly 10 mCrab. All of these latest detections show the 401-Hz pulsation, and some of them also again show violent x-ray flaring. Given the highly erratic behavior of the x-ray source, we strongly encourage monitoring observations at all wavelengths.”

XTE J1543–568 T. Takeshima, Goddard Space Flight Center (GSFC) and Universi-

ties Space Research Association (USRA); F. E. Marshall, GSFC; R. H. D. Corbet, GSFC and USRA; and J. in ’t Zand, Space Research Organization Netherlands, report an improved position for the 27.12-s pulsar XTE J1543–568 (IAUC 7363) from an RXTE/PCA observation on Feb. 14.04 UT: “The best-fit position is α = 15^{h}44^{m}13^{s}, δ = −56^{o}44′.5 (equinox 2000.0; 90-percent confidence error radius 4′). The source flux and the pulsed fraction remain at ∼ 6 mCrab and 60 percent (2–10 keV), respectively. Observations at other wavelengths are encouraged.”

CORRIGENDA On IAUC 7158, ‘Supernovae 1999ac, 1999ad, 1999an, 1999be’, line 7,

for SN 1999ad read SN 1999ac

On IAUC 7314, ‘Supernova 1999fz in UGC 8164’, lines 5–6, for 2′′.0 west read 5′′.9 west

On IAUC 7358, ‘Supernova 2000G in UGC 1773’, line 5, for 0′′.13 east read 1′′.9 east

2000 February 28  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green